# Unsupervised machine learning for the detection of exotic phases in skyrmion phase diagrams

### v1

@@ IV.1 Application to ferromagnetic skyrmion models
The KFDMz model has been shown to have several exotic behaviors, due to the competition between skyrmion and chiral spin liquid physics [35, 49]. Besides the “bimeron glass” above mentioned, for the value of chosen here, it has been shown that, coming from the high temperature paramagnetic phase, the skyrmion lattice is formed by the discrete emergence of skyrmions at intermediate temperatures, forming a skyrmion gas that gets more densely populated as the temperature is lowered. We compare the RMSE and BCE as a function of temperature for the two ferromagnetic kagome skyrmion models, at three different values of the magnetic field . In panels (a) and (b) it can be seen that for the KFDMxy model both RMSE and CAE go down smoothly with temperature. The BCE is smaller for higher fields, but the RMSE is similar for both . This is simply due to the fact that for this model at those magnetic fields the system is in a skyrmion crystal phase (see Fig. 3, left panel). The variables hint a different physical process for the KFDMz model. At , both the RMSE and BCE abruptly go up at intermediate temperatures, a feature which is much sharper in the BCE curve. As discussed in previous works [35, 49], for this the KFDMz model goes through an skyrmion gas phase at intermediate temperatures, with a field polarized chiral spin liquid background. Since this type of phase is better decoded by the CAE than the bimeron glass phase, the BCE and RMSE are smaller. This type of intermediate phase is also present at , which is seen as a small well before the “bump” in the BCE curve. We show an example MC snapshot in this phase and its decoded image in Fig. 4, bottom row. Since the temperature is higher and the background is a polarized chiral spin liquid, it can be seen that it is not completely aligned with the field (i. e. the bakcround is not completely red). On the other hand, this changes when decoding the image with the CAE, where thermal fluctuations are smoothed. This is related to the potential “denoising” use of convolutional autoencoders, showing that these algorithms, if needed, may also be considered as tools to “erase” fluctuations and better define structures such as skyrmions.
@@
As a summary, we present a comparison of the phase diagrams of the (third nearest neighbor) chirality, RMSE and BCE from CAE for the two kagome models, in Fig. 7. For both models, we first see that the RMSE and BCE are higher at higher temperature, as expected. Then, for BCE, at the lower temperatures it goes down with magnetic field and is lowest when the system is in the field polarized phase, and it does not seem to easily distinguish the lower field phases. However, at intermediate temperatures we see a particular behaviour for the KFDMz case: a drop in the BCE, which matches the chiral spin liquid region [35, 49]. As for the RMSE, in the KFDMxy model at lower temperatures it distinguishes between three regions, with a higher value at lower magnetic field (corresponding to helices and bimerons), lowest values at intermediate magnetic fields (skyrmions phase) and then goes a bit up at higher fields, in the field polarized region. An important difference is observed for KFDMz: a small region at higher field where the RMSE goes up again, between the skyrmion lattice phase and the field polarized one. This is another exotic feature of this model, a higher field bimeron phase with potentially glassy characteristics [35], which is not evident inspecting the chirality. Thus, we see that the RMSE, and also BCE, are powerful tools to pinpoint regions in parameter space where a system may deviate from the “typical” skyrmion phase diagram. Given that we only analyse snapshots, this technique may be applied to real-space images, and provide helpful insight of a skyrmion system without resorting to designing a model and permorfing simulations.
@@ IV.2 Application to antiferromagnetic models
Although the CAE was trained on ferromagnetic triple-q skyrmion lattices, in this subsection we aim to explore whether this CAE may give some insight in two well-known antiferromagnetic models in the triangular lattice, presented in Table 2. For these models, the exchange interaction is antiferromagnetic (), which implies a frustrated system due to the lattice geometry. In the first model, TAFDMxy, there are in-plane DMI (). It has been shown that under a magnetic field, at finite temperature an antiferromagnetic skyrmion lattice formed by three interpenetrated sublattices is stabilized [36], a texture that has been shown to be stabilized in other frustrated models [40, 58, 59], and connected to the fractional antiferromagnetic skyrmions in MnSc2S4 [37, 60]. As a second model we consider the pure exchange antiferromagnetic model (), where the Hamiltonian may be rewritten as a sum of the total spin of triangular plaquettes. Under a magnetic field , a notable feature of this model is that at at finite temperature order-by-disorder (entropic selection) induces a pseudo-plateau at magnetization , where in each plaquette two spins are parallel to the external field (“up”, ) and one is antiparallel (“down”, ), and is thus known as a pseudo-plateau [61, 62, 63]. For lower fields, there is a coplanar “Y” state, and for higher fields a coplanar “V” state [61].
@@
Clearly, the emergent antiferromagnetic low-temperature textures from these models are quite different from the triple-q ferromagnetic skyrmion lattices, and thus a significantly larger RMSE is expected. To explore whether some information on these models may nonetheless be extracted with the skyrmion-trained CAE, we proceed as before: we apply the CAE to low temperature snapshots obtained from MC simulations of both models, and calculate the RMSE and the BCE.
@@
In Fig. 8 (a) and (b) we show the RMSE, BCE and chirality as a function of the external field for the TAFDMxy and TAF models, respectively, at the lowest simulated temperature ). In the TAFDMxy case (panel (a)), the RMSE and the BCE are relatively flat at lower and present a sharp decrease at , which matches the jump in the chirality associated with the stabilization of the AF skyrmion lattice. Then as is increased both quantities go down, and the curves get steeper when drops to zero. Although the behaviour of the RMSE and the BCE is similar, the features in the RMSE are sharper than for BCE, particularly the steep drop at lower field. So, in this case, the RMSE shows that there is a first type of phase stabilized at , followed by a different type of phase up to , where the RMSE starts to drop until it flattens at its minimum value at higher fields, associated to the ferromagnetic phase, with all spins aligned with .
@@
The BCE and RMSE curves are quite different for the TAF case (panel (b)): the BCE goes down with increasing magnetic field, presenting a small kink at , the pseudo-plateau magnetic field. On the other hand, the RMSE curve has two clearly distinct regions, separated at : a low magnetic field region where the RMSE goes up, and a higher field region where it goes down, presenting thus a peak at . Therefore, even if, as expected since the CAE was trained for skyrmion lattices, the obtained RMSE values are quite large, the RMSE calculated after applying this CAE to the MC snapshots indicates that there are two types of phases at lower and higher field, and that there is a particular feature in the TAF model at .
@@
Panels (c) and (d) show the RMSE and BCE curves, respectively, for these two AF models and the simple ferromagnetic skyrmion model (TFDMxy; here the magnetic field has been rescaled for a better comparison of the magnitude and characteristics of the curve). As expected, the magnitude of both quantities, particularly RMSE, is significantly larger for the antiferromagnetic models. The three RMSE curves, contrary to the BCE ones, have different characteristics, showing that the three models have different types of phases. The kinks and jumps in the three curves are also indicators of the different phases within each model.
@@ V Conclusions
Secondly, we applied the CAE to two models in the kagome lattice, the well known skyrmion model we had simulated in the triangular lattice, and one where an additional out-of-plane DMI is included. As expected, the behaviour of the RMSE in the first case (only planar DMI) is similar to that in the triangular lattice. However, in the second case (with both planar and out-of-plane DMI) there is a clear region at low temperature and higher fields where the error raises, indicating the possibility of a different type of phase. In fact, previous studies in this model [35] show that there is a high field bimeron glass phase. Moreover, this model also has a variety of phases at higher temperature, which is reflected in both the RMSE and BCE. Our analysis also suggests that the spread of the RMSE, considering independent realizations for the MC data, may also be a tool to distinguish between different low temperature phases.
@@
Finally, we apply the skyrmion-lattice trained CAE to snapshots from two models in the antiferromagnetic triangular lattice: one with in-plane DMI, where antiferromagnetic skyrmion lattices are stabilized, and the pure exchange model, which presents a well known plateau at . Given that the CAE is trained with ferromagnetic skyrmion lattices, a good decodification of the input data is not expected. Nonetheless, the low-temperature RMSE curves have features that suggest possible different phases. In particular, in the pure antiferromagnetic exchange model, the RMSE is maximum for values of the magnetic field where the pseudo-plateau emerges.
@@
In conclusion, we have shown how the anomaly detection technique, which we have applied resorting to a very simple Convolutional Autoencoder trained with analytically generated skyrmion lattices and choosing the RMSE to measure the error between input and output data, may give relevant information on skyrmion systems. We have demonstrated that it was able to distinguish the three typical phases in a skyrmion phase diagram (helices, skyrmions and field-polarized spins), and to spot regions in parameter space where there may be exotic behavior. Additionally, we have seen how it may also hint the existence of different phases in other types of models. Thus, we expect our work to support and further promote the exploration and use of machine learning techniques in different magnetic models.

### v2

@@ IV.1 Application to ferromagnetic skyrmion models
The KFDMz model has been shown to have several exotic behaviors, due to the competition between skyrmion and chiral spin liquid physics [40, 59]. Besides the “bimeron glass” above mentioned, for the value of chosen here, it has been shown that, coming from the high temperature paramagnetic phase, the skyrmion lattice is formed by the discrete emergence of skyrmions at intermediate temperatures, forming a skyrmion gas with well formed skyrmions in a field polarized background with chiral spin liquid properties that gets more densely populated as the temperature is lowered. To explore whether this analysis can shed light on this phase, in Fig. 8 we compare the RMSE and BCE as a function of temperature for the two ferromagnetic kagome skyrmion models, at three different values of the magnetic field . In panels (a) and (b) it can be seen that for the KFDMxy model both RMSE and BCE go smoothly down to a given value as the temperature is lowered. This given value for BCE gets smaller as the magnetic field is increased (panel (b)), but for the RMSE (panel (a)) the low temperature value is similar for both . This is simply due to the fact that for this model at those magnetic fields the system is in a skyrmion crystal phase (see Fig. 5, left panel). The variables hint a different physical process with temperature for the KFDMz model, Fig. 8, panels (c) and (d). Both the RMSE and BCE curves show a sharp feature at , coming from a higher value at higher temperatures (), they go down with temperature, show a minimum at intermediate temperatures () and then jump to settle to their low temperature value. A similar feature is seen for the BCE at .
As discussed in previous works [40, 59], for these the KFDMz model goes through an skyrmion gas phase at intermediate temperatures, with a field polarized chiral spin liquid background. Clearly this type of phase is better decoded by the CAE than the bimeron glass phase found at lower temperatures, and therefore both the BCE and RMSE show a minimun at intermediate temperature. We present an example MC snapshot of this phase and its decoded image in Fig. 9. Since the temperature is higher and the background is a field polarized chiral spin liquid, it can be seen that it is not completely aligned with the field (i. e. the background is not completely red). On the other hand, this changes when decoding the image with the CAE, where thermal fluctuations are smoothed. This is related to the potential “denoising” use of convolutional autoencoders, showing that these algorithms, if needed, may also be considered as tools to “erase” fluctuations and better define structures such as skyrmions.
@@
As a summary, we present a comparison of the phase diagrams of the chirality density obtained from MC simulations, and the calculated RMSE and BCE from CAE for the two kagome models, in Fig. 10. For both models, we first see that the RMSE and BCE are higher at higher temperature, as expected. Then, for BCE, at the lower temperatures it goes down with magnetic field and is lowest when the system is in the field polarized phase, and it does not seem to easily distinguish the lower field phases. However, at intermediate temperatures we see a particular behaviour in the KFDMz case: a drop in the BCE, which matches the chiral spin liquid region [40, 59]. As for the RMSE, in the KFDMxy model at lower temperatures it distinguishes between three regions, with a higher value at lower magnetic field (corresponding to helices and bimerons), lowest values at intermediate magnetic fields (skyrmions phase) and then goes a bit up at higher fields, in the field polarized region. An important difference is observed for KFDMz: a small region at higher field where the RMSE goes up again, between the skyrmion lattice phase and the field polarized one. This is another exotic feature of this model, a higher field bimeron phase with potentially glassy characteristics [40], which is not evident inspecting the chirality. Thus, we see that the RMSE, and also BCE, are powerful tools to pinpoint regions in parameter space where a system may deviate from the “typical” skyrmion phase diagram. Given that we only analyse snapshots, this technique may be applied to real-space images, and provide helpful insight of a skyrmion system without resorting to designing a model and permorfing simulations.
@@ IV.2 Application to antiferromagnetic models
Although the CAE was trained on ferromagnetic skyrmion lattices, in this subsection we aim to explore whether this CAE may give some insight in two well-known antiferromagnetic models in the triangular lattice, presented in Table 2. For these models, the exchange interaction is antiferromagnetic (), which implies a frustrated system due to the lattice geometry. In the first model, TAFDMxy, there are in-plane DMI (). It has been shown that under a magnetic field, at finite temperature an antiferromagnetic skyrmion lattice formed by three interpenetrated sublattices is stabilized [43], a texture that has been shown to be stabilized in other frustrated models [47, 70, 71], and connected to the fractional antiferromagnetic skyrmions in MnSc2S4 [44, 72]. As a second model we consider the pure exchange antiferromagnetic model (), where the Hamiltonian may be rewritten as a sum of the total spin of triangular plaquettes. Under a magnetic field , a notable feature of this model is that at at finite temperature order-by-disorder (entropic selection) induces a pseudo-plateau at magnetization , where in each plaquette two spins are parallel to the external field (“up”, ) and one is antiparallel (“down”, ), and is thus known as a pseudo-plateau [73, 74, 75]. For lower fields, there is a coplanar “Y” state, and for higher fields a coplanar “V” state [73].
@@
Clearly, the emergent antiferromagnetic low-temperature textures from these models are quite different from the ferromagnetic skyrmion lattices, and thus a significantly larger RMSE is expected. To explore whether some information on these models may nonetheless be extracted with the skyrmion-trained CAE, we proceed as before: we apply the CAE to low temperature snapshots obtained from MC simulations of both models, and calculate the RMSE and the BCE.
@@
In Fig. 11 (a) and (b) we show the RMSE, BCE and chirality as a function of the external field for the TAFDMxy and TAF models, respectively, at the lowest simulated temperature ). In the TAFDMxy case (panel (a)), the chirality curve is similar to that from the TFDMxy models in Fig. 4, and therefore one might expect similar low temperature magnetic textures from this MC order parameter. Inspecting the RMSE and BCE curves, there is a strong hint that something different is going on, since the errors are higher and the curves as a function of are not the same. The RMSE and the BCE are relatively flat at lower and present a sharp decrease at , which matches the jump in the chirality associated with the stabilization of the AF skyrmion lattice. Then as is increased both quantities go down, and the curves get steeper when drops to zero. Although the behaviour of the RMSE and the BCE is similar, the features in the RMSE are sharper than for BCE, particularly the steep drop at lower field. So, in this case, the RMSE shows that there is a first type of phase stabilized at , followed by a different type of phase up to , where the RMSE starts to drop until it flattens at its minimum value at higher fields, associated to the ferromagnetic phase, with all spins aligned with . Moreover, since it is not at a minimum where the chirality is highest, it also suggests that the textures in these range of magnetic field are not ferromagnetic skyrmion lattices, something that is not obvious by simply inspecting the chirality curve from MC.
@@
The BCE and RMSE curves are quite different for the TAF case (panel (b)): the BCE goes down with increasing magnetic field, presenting a small kink at , the pseudo-plateau magnetic field. On the other hand, the RMSE curve has two clearly distinct regions, separated at : a low magnetic field region where the RMSE goes up, and a higher field region where it goes down, presenting thus a peak at . Therefore, even if, as expected since the CAE was trained for skyrmion lattices, the obtained RMSE values are quite large, the RMSE calculated after applying this CAE to the MC snapshots indicates that there are two types of phases at lower and higher field, and that there is a particular feature in the TAF model at , since the error here is highest and the curve is sharp.
@@
Panels (c) and (d) show the RMSE and BCE curves, respectively, for these two AF models and the simple ferromagnetic skyrmion model TFDMxy; here the magnetic field has been (scaled to its saturation value in each model, ) for a better comparison of the magnitude and characteristics of the curve . As expected, the magnitude of both quantities, particularly RMSE, is significantly larger for the antiferromagnetic models. The three RMSE curves, contrary to the BCE ones, have different characteristics, showing that the three models have different types of phases. The kinks and jumps in the three curves are also indicators of the different phases within each model.
@@
IV.3 Application to the square lattice
Throughout this work, we have focused on triangular based lattices. We now explore the portability of the trained CAE, applying it to snapshots from simulations for the Hamiltonian in Eq.(3) in the square lattice for in-plane DMI (SFDMxy). As in the previous cases, in Fig. 13 we plot, for the lowest simulated temperature (), the RMSE and BCE between the input and output images as a function of the external magnetic field, and we also show the chirality density calculated from simulations to analyze what information may be extracted from the behavior of the RMSE and BCE.
As expected, the behavior of the chirality in this lattice is as in the triangular lattice case: as a function of the magnetic field, there is an intermediate region of non-zero chirality, related to the formation of skyrmions. Surprisingly, we see that, even if we applied a CAE trained with triangular based lattices, both the BCE and the RMSE follow a trend similar to that seen in the TFDMxy models in Fig. 4. Most importantly, the RMSE is also minimum for the skyrmion phase, with a value comparable to the triangular lattice models. This may not be expected, since the CAE reads the input data as a matrix, and thus a skyrmions in a triangular lattice are actually seen more elongated. In Fig. 14 we show two examples of input and output configurations for the square lattice, corresponding to a helical and a skyrmion lattice configuration, showing that the outputs retain the main features of these phases even in a different lattice geometry.
@@ V Conclusions
Secondly, we applied the CAE to two models in the kagome lattice, the well known skyrmion model we had simulated in the triangular lattice, and one where an additional out-of-plane DMI is included. In both cases, the chirality from MC simulations, which is the order parameter for skyrmions, is quite similar. Regarding the errors, as expected, the behaviour of the RMSE in the first case (only planar DMI) is similar to that in the triangular lattice. However, in the second case (with both planar and out-of-plane DMI) there is a clear region at low temperature and higher fields where the error raises, indicating the possibility of a different type of phase(an “anomaly”). In fact, previous studies in this model [40] show that there is a high field bimeron glass phase. Moreover, this model also has a variety of phases at higher temperature, which is reflected in both the RMSE and BCE. Our analysis also suggests that the spread of the RMSE, considering independent realizations for the MC data, may also be a tool to distinguish between different low temperature phases.
@@
Moreover, we apply the skyrmion-lattice trained CAE to snapshots from two models in the antiferromagnetic triangular lattice: one with in-plane DMI, where antiferromagnetic skyrmion lattices are stabilized, and the pure exchange model, which presents a well known plateau at . Given that the CAE is trained with ferromagnetic skyrmion lattices, a good decodification of the input data is not expected. Nonetheless, the low-temperature RMSE curves have features that suggest possible different phases. For the antiferromagnetic skyrmion model, the chirality curve does not indicate significant differences from the ferromagnetic case. However, the RMSE curves do, and the values of the RMSE are higher than in the ferromagnetic case, signaling that the resulting textures with non zero-chirality are not ferromagnetic skyrmion lattices. Then, in the pure antiferromagnetic exchange model, the RMSE is maximum for values of the magnetic field where the pseudo-plateau emerges.
Finally, we explore the possibility of applying the CAE trained with triangular lattices in other geometries. To do this, we repeated the analysis for a ferromagnetic skyrmion model in the square lattice, and showed that the RMSE from CAE is able to distinguish between the three main low temperature phases, thus expanding the possibilities of application.
@@
In conclusion, we have shown how the anomaly detection technique, which we have applied resorting to a very simple Convolutional Autoencoder trained with analytically generated skyrmion lattices and choosing the RMSE to measure the error between input and output data, may give relevant information on skyrmion systems, which may not be easily determined simply by studying the chirality parameter. We have demonstrated that the comparison and study of the RMSE was able to distinguish the three typical phases in a skyrmion phase diagram (helices, skyrmions and field-polarized spins), and to spot regions in parameter space where there may be exotic behavior. Additionally, we have seen how it may also hint the existence of different phases in other types of models, and that it may be applied to skyrmion models in other lattice geometries. Thus, we expect our work to support and further promote the exploration and use of machine learning techniques in different magnetic models. On the one hand, these techniques complement the analysis from simulations, and, on the other, since snapshots are treated as images, these techniques may even be applied to experimental data obtained with spin-polarized scanning tunneling microscopy techniques [42, 64, 65, 66].
@@ Acknowledgments
Appendix
Here we discuss the effect of changing parameters of the PCA and CAE in our analysis. First, we focus on the PCA. Aiming for more variability, we reduce the number of principal components of the PCA from to in order to reach an accumulated variance of (PCA-2). Even if this may imply a larger RMSE, it could also imply larger differences between the helices and the skyrmion phases, for example, since it would retain the main features of skyrmion lattices. However, we find that this is not the case, we present the PCA-2 RMSE as a function of at the lowest simulated temperature for TFDMxy in Fig. 15, to be contrasted with the upper panel of Fig. 4. Although the RMSE seems to be a bit larger in the helical phase, the differences between the helical and the skyrmion region are be enough to distinguish between the two phases within the errors.
Secondly, we construct a different CAE. The CAE used in the main work is very simple and only has one convolutional and one MaxPooling layer for the enconding part. We construct a deeper CAE (CAE-2) adding in the encoder an extra convolutional layer with 64 filters of size 4 and MaxPooling layer with filter size 3, and thus in the decoder we add a convolutional and an UpSampling layer. The rest of the structure and the training method remains unchanged. We get a RMSE of 0.0425 in the training set, and 0.0432 in the validation set. To explore whether a change in the CAE affects the main result of this work, i. e. the possibility to distinguish phases in a skyrmion phase diagram and to detect regions with possible exotic textures that may not be easily distinguished with MC order parameters, we apply the CAE-2 to the ferromagnetic models. Results for TFDMxy are shown in Fig. 15: the RMSE is clearly different for the helices and the skyrmion phases, but the distinction is not that clear between the skyrmion phases and the ferromagnetic order. There is a single point with a slightly larger value, that corresponds to the skyrmion gas phase, which is narrow at lower temperatures. The BCE retains its overall behavior (it is highest in the helices and smallest in the ferromagnetic case), although the jumps between the three main phases are sharper.
To explore whether CAE-2 may detect exotic phases, we compare the results of the RMSE and BCE phase diagrams for the kagome model KFDMz in Fig. 16, to contrast with Fig. 10. We find that, in general, the RMSE is bigger, implying that the reconstruction is not as good as in the first proposed CAE. However, different regions may more easily be distinguished in temperature and magnetic field, and the anomaly detection technique comparing the RMSE in different regions may still work. Most interestingly, the exotic phases (the high field bimeron glass and the intermediate skyrmion gas phases) correspond to larger changes both in the RMSE and in the BCE phase diagrams. However, the larger RMSE values are related to poorer reconstructions of the spin configurations, as we show below. In Fig. 17 we present the MC snapshot and decoded images for helices and skyrmion lattice for the TFDMxy model (to be compared with Fig. 3) and for the bimeron glass and the skyrmion gas that emerge in the KFDMz case (to be compared with Figs. 6,9). Clearly, the CAE-2 fails to reconstruct features that are not skyrmions, and even in the skyrmion lattice case the decoded configuration looses definition when compared to the first CAE model, giving the impression of “blurred” images. Although a good reconstruction is not the key of the technique, the lack of significant differences in the RMSE in the skyrmion and ferromagnetic regions could imply that these types of phases may not be distinguished in other contexts. .
Finally, we apply the CAE-2 to the ferromagnetic square lattice model. Contrary to the ferromagnetic cases in triangular based lattices discussed above, we find that, although the BCE curve is similar to that of the TFDMxy model presented in Fig. 15, here the RMSE in the skyrmion phase is higher than in the ferromagnetic phase, as shown in the top panel of Fig. 18 where we plot the BCE and the RMSE from CAE-2 as a function of magnetic field at the lowest simulated temperature, to be compared with Fig. 13. In the bottom panel of Fig. 18 we present examples of input and output snapshots obtained with CAE-2 for a helical and a skyrmion lattice configuration in the square lattice. It can be seen that CAE-2 is more adapted to the triangular based lattices, since in the square lattice it “elongates” skyrmions, and therefore the decoded images are significantly different to those presented in Fig. 14.
In conclusion, a deeper CAE would be more specialized in the specific skyrmion lattices used for training. On the one hand, this could be desirable to distinguish other configurations from the RMSE, specially with the intention to detect exotic orderings. On the other hand, in this case, compared to the first CAE, it implied a poorer decodification for skyrmions in other geometries (the square lattice), and even for skyrmions in triangular lattices, which may not match exactly those constructed from the analytical parametrisation (Eq.(1)) used for training and validation of the algorithms, reducing the difference in the RMSE between the skyrmion and ferromagnetic phases, which is the key in the anomaly detection techinique.